# Stackelberg vs. Nash in the Lottery Colonel Blotto Game

### v1

@@ Abstract
Resource competition problems are often modeled using Colonel Blotto games. However, Colonel Blotto games only simulate scenarios where players act simultaneously. In many real-life scenarios, competition is sequential, such as cybersecurity, cloud services, business investments, and more. To model such sequential competition, we model the Lottery Colonel Blotto game as a Stackelberg game. We solve the Stackelberg equilibrium in the Lottery Colonel Blotto game in which the first mover’s strategy is actually a solution to a bi-level optimization problem. We develop a constructive method that allows for a series of game reductions. This method enables us to compute the leader’s optimal commitment strategy in a polynomial number of iterations. Furthermore, we identify the conditions under which the Stackelberg equilibrium aligns with the Nash equilibria. Finally, we show that by making the optimal first move, the leader’s utility can be unbounded compared to its utility in Nash equilibria. We find that the player with a smaller budget has a greater incentive to become the leader in the game. Surprisingly, even when the leader adopts the optimal commitment strategy, the follower’s utility may improve compared to that in Nash equilibria.
@@ 1 Introduction
Competing for resources with limited budgets in strategic settings has a wide range of applications. For example:
Advertising auctions and e-Commerce markets. Advertisers compete for limited ad placements on online platforms, allocating their budgets to maximize exposure and conversion rates. Large enterprises often act as leaders, with smaller retailers responding as followers. These budget strategies significantly impact the ad visibility and conversions.
Economic aspects of Web, GenAI, and cloud computing. Providers like AWS and Google Cloud compete across multiple markets (e.g., geographic regions or industries), allocating cloud resources to attract customers and maximize market share. One provider acts as the leader, allocating resources first to focus on high-value markets. The other, as the follower, observes the leader’s strategy and optimizes its own allocation, targeting underdeveloped or less saturated markets.
@@
These examples, along with many others from areas such as electoral competition (Behnezhad et al., 2018), security (Pagey et al., 2023; Zhao et al., 2023; Ferdowsi et al., 2017; Gupta et al., 2014), crowdsourcing (Roitero et al., 2022; Haggiag et al., 2022; AlShehry and Ferguson, 2015), and recommendation systems (Anwar et al., 2024; Lee et al., 2022), are modeled by Colonel Blotto games (Borel, 1921). These types of games have attracted considerable interest from sociology (Xiao et al., 2020; Fu and Iyer, 2019), economics (Jackson and Nei, 2015), and computer science (Kohli et al., 2012).
@@
There are many variants of Colonel Blotto games, in which players simultaneously allocate their limited budgets (resources) across various battlefields. These budgets can be represented by integers, such as troops, or by real numbers, such as bid prices in auctions. The player’s utility on a single battlefield is determined by the resources invested by players in that battlefield and the outcome function of that battlefield. A player’s payoff is then calculated by aggregating the utility gained from all battlefields. Based on the outcome function of the battlefield, Colonel Blotto games can be categorized into two types: one where the outcome function follows the winner-takes-all rule, and the other where the outcome function is proportional. Colonel Blotto games with a proportional outcome function are also referred to as the Lottery Colonel Blotto game. In this paper, we focus on the Lottery Colonel Blotto game.
@@
Although Colonel Blotto games are highly versatile, most existing research treats them as normal-form games, primarily focusing on equilibrium existence, computation, and closed-form characterizations in restricted settings (Roberson, 2006; Hortala-Vallve and Llorente-Saguer, 2012; Macdonell and Mastronardi, 2015; Perchet et al., 2022). However, in many real-world situations, such as the Web, cloud computing, cybersecurity, and advertising auction examples, one player can commit to a strategy before the other player makes their decision.
In this paper, we model the Lottery Colonel Blotto game as a Stackelberg game. The Stackelberg game describes a strategic scenario in which one player, as the leader (first mover), determines strategy, while the other player, as the follower (second mover), responds based on the leader’s strategy (Hicks, 1935; Schelling, 1960).
@@
The Stackelberg game can be viewed as a sequential game. In many previous studies, the focus has been on scenarios involving an attacker, as a follower, and a defender, as a leader, both operating under cost constraints rather than budget limitations. Under these cost constraints, they determine their respective efforts in order to maximize their individual payoffs (Zhuang and Bier, 2007; Cavusoglu et al., 2008; Hausken et al., 2008; Hausken, 2012). However, Iliaev et al. (2023) point out that the model without budget constraints is easier to solve analytically because it is equivalent to solving a single-battlefield model with $n$ different values. Iliaev et al. (2023) consider budget constraints to compare the sequential and the simultaneous game. However, they only impose budget constraints on the follower, while the first-mover still has only costs without budget limitations, and both sides have symmetric valuations for the same battlefield. Here, for broader applicability, we consider a more general model in which both the leader and the follower have budgets. In addition, both players and the battlefield values are asymmetric. Furthermore, we will compare the sequential game and the simultaneous game, i.e., the Stackelberg equilibrium vs. Nash equilibria.
In addition to the above dynamics, there exists another type of dynamics. All players simultaneously allocate resources in the first battlefield, then allocate resources in the second battlefield, and so on (Xie and Zheng, 2022; Anbarci et al., 2023).
@@
To compare sequential games and simultaneous games, we need to analyze Stackelberg equilibrium and Nash equilibria. However, finding a closed-form representation of Nash equilibria in Colonel Blotto games is a known challenge (Perchet et al., 2022; Li and Zheng, 2022). Such representations have been characterized only in specific scenarios, such as when battlefield values are identical (Roberson, 2006) or when players have symmetrical budgets (Hortala-Vallve and Llorente-Saguer, 2012; Boix-Adserà et al., 2020). Most existing research offers only partial characterizations, focusing on conditions for the existence and uniqueness of Nash equilibria. For the Stackelberg equilibrium, we need to compute the leader’s optimal commitment. However, previous work has provided negative conclusions. Finding an optimal pure strategy to commit to is NP-hard in the normal-form games (Conitzer and Sandholm, 2006; Korzhyk et al., 2010; Letchford and Conitzer, 2010), as it requires solving a bi-level optimization problem (Renou, 2009). To address this, we investigate new insights into optimal commitment and best-response dynamics, enabling a comparison between Stackelberg equilibrium and Nash equilibria in these games.
@@ 1.1 Our Contribution
  • By reformulating the leader’s objective and reducing the search space for the follower’s best response strategy, we can characterize the leader’s optimal commitment strategy through a polynomial number of iterations.
  • We characterize the sufficient and necessary conditions, under which the Stackelberg equilibrium coincides with the Nash equilibria of the simultaneous-move game.
  • Finally, we address our motivating questions by constructing extreme cases that compare the leader’s utility when using the optimal commitment strategy to their utility in Nash equilibria. We observe that this ratio is also dependent on the leader’s budget relative to the follower’s budget. Furthermore, we provide an example that illustrates how the follower’s utility may still increase when faced with the leader’s optimal commitment.
@@ 1.2 Related Work
When the outcome function is winner-takes-all, the probability of a player winning a battlefield depends on the proportion of resources they allocate compared to the total resources allocated by all players. This type of Colonel Blotto games is also called the Lottery Colonel Blotto game. In this line of work, it is called symmetric if all players have equal budgets; otherwise, it is asymmetric. When all battlefields have the same value, and this value is consistent for all players, the game is homogeneous. If the battlefields possess different values, but these values remain the same for every player, the game is heterogeneous. Finally, when the battlefields’ values differ, and these values vary for each player, the game is termed a generalized game. The Lottery Colonel Blotto game has been proven to possess pure Nash equilibria (Kim et al., 2018). Friedman (1958) shows that the pure Nash equilibrium is unique and exhibits proportionality features in a two-player setup with heterogeneous battlefield values and asymmetric player budgets. Duffy and Matros (2015) extend Friedman’s analysis from two players to multiple players. Kim et al. (2018) provide a method to identify all pure Nash equilibria in the case of two players with asymmetric budgets and multiple battlefields with generalized values, and present an example where the pure Nash equilibrium is not unique. Kovenock and Arjona (2019) provide the best response given the strategy of the other player. However, none of the existing work has provided a closed-form representation of the pure Nash equilibria (Kim et al., 2018; Kovenock and Arjona, 2019). In a variant where the outcome function of a battlefield is determined by the Tullock rent-seeking contest success function, the proportional rule is parameterized by a battle-specific discriminatory power and a battle-and-contestant-specific lobbying effectiveness (Xu and Zhou, 2018; Osorio, 2013; Li and Zheng, 2022). In this variant, the existence of a pure Nash equilibrium is not always guaranteed. Xu et al. (2022) provide sufficient conditions for the existence of the pure Nash equilibria, while Li and Zheng (2022) analyze the properties of the pure Nash equilibria and provide sufficient conditions for its uniqueness.
@@ 2 Preliminaries
$u_{i}(\bm{x}_{a},\bm{x}_{b})=\sum_{j=1}^{n}u_{ij}=\sum_{j=1}^{n}\frac{x_{ij}}{x_{ij}+x_{-ij}}\cdot v_{ij},\quad\forall i\in\{a,b\}.$ (2)
When $x_{aj}>0$, for all $j\in[n]$, Kovenock and Arjona (2019) characterize the other player’s best response functions as follows. This characterization is a useful tool for our subsequent analysis. To use this result to characterize the leader’s optimal commitment, we first assume that the leader will allocate a positive budget to every battlefield. Following a series of constructions in Section 3, we identify the support of the follower’s best response strategies and provide a closed-form expression of the leader’s optimal commitment strategy in Section 4. To complete this analysis, we will verify that the leader indeed allocates a positive budget to every battlefield in the optimal commitment strategy. We focus on the pure strategy of the leader.
@@
(Kovenock and Arjona, 2019) Given the leader’s strategy $\bm{x}_{a}=(x_{aj})_{j=1}^{n}$, where $x_{aj}>0$, $\forall j\in[n]$, assume without loss of generality that all battlefields are ordered such that $\frac{v_{b1}}{x_{a1}}\geq\frac{v_{b2}}{x_{a2}}\geq\cdots\geq\frac{v_{bn}}{x_{an}}$. The unique optimal budget allocation of the follower, $\bm{x}_{b}=(x_{b1},x_{b2},\cdots,x_{bn})$, to battlefield $j\in[n]$ is characterized as follows:
$x_{bj}=\begin{cases}\frac{\left(x_{aj}v_{bj}\right)^{\frac{1}{2}}\cdot\left(x_{b}+\sum_{j^{\prime}\in K(\bm{x}_{a})}x_{aj^{\prime}}\right)}{\sum_{j^{\prime}\in K(\bm{x}_{a})}\left(x_{aj^{\prime}}v_{bj^{\prime}}\right)^{\frac{1}{2}}}-x_{aj},&\text{ if }j\in K(\bm{x}_{a});\\ 0,&\text{ if }j\in[n]\backslash K(\bm{x}_{a}),\end{cases}$ (3)
@@
$k^{*}=\max\left\{k\in[n]:\frac{v_{bj}}{x_{aj}}>\frac{\left(\sum_{l=1}^{j}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum_{l=1}^{j}x_{al}\right)^{2}},\quad\forall j\leq k\right\}.$ (4)
@@
Due to space limitations, most proofs are deferred to the Appendix. However, we provide technical remarks and insights into the proofs within the main text.
@@ 3 The Support of the Follower’s Best Response Strategy
    Consider a game $\mathcal{G}$, with the leader’s commitment strategy denoted by $\bm{x}_{a}$ (which may not be optimal), and the follower’s best response denoted by $\bm{x}_{b}$. If we divide a single battlefield into multiple sub-battlefields in such a way that the leader and follower’s valuations of the original battlefield are evenly distributed among these sub-battlefields, the resulting game is denoted by $\mathcal{G}^{(1)}$. By evenly distributing their budgets across these sub-battlefields, the utilities for both the leader and the follower remain unchanged in $\mathcal{G}^{(1)}$. The formal statement and its proof are deferred to Lemma 7 in the Appendix A due to space constraints.
@@
    By repeatedly performing this battlefield-splitting process, we create a game in which the follower values all battlefields equally. Without loss of generality, we can rename the battlefields so that the leader values them in increasing order. Denote this game by $\mathcal{G}^{(2)}$. According to Lemma 8 in the Appendix A, in the optimal commitment strategy, the leader should allocate more or at least equal budget to battlefields that they value strictly higher in order to achieve a strictly greater payoff.
@@
    Furthermore, we observe that, in the optimal commitment strategy, when the leader values two battlefields equally, they should allocate an equal budget to both battlefields. For a formal statement and proof, see Lemma 9 in the Appendix A.
@@
    Following the above procedure, we obtain a game where $v_{b1}=v_{b2}=\dots$, meaning the follower values all battlefields equally. Meanwhile, the leader values the battlefields in increasing order, with some values strictly increasing and others remaining equal. Our final operation is to merge these battlefields that split from the same battlefield. In this new game $\mathcal{G}^{(3)}$, the leader and follower’s values and budgets across these sub-battlefields are merged into the original battlefield. Their utilities remain unchanged from $\mathcal{G}^{(2)}$. The formal statement and proof are deferred to Lemma 10 in the Appendix A.
With these constructions and observations, we can now prove the main result of this section.
@@
$\frac{\partial u_{bj}}{\partial x_{bj}}=\frac{\partial(\frac{x_{bj}v_{bj}}{x_{aj}+x_{bj}})}{\partial x_{bj}}=\frac{x_{aj}v_{bj}}{(x_{aj}+x_{bj})^{2}},\quad\frac{\partial^{2}u_{bj}}{\partial x_{bj}^{2}}=\frac{-2x_{aj}v_{bj}}{(x_{aj}+x_{bj})^{3}}<0.$ (5)
Hence, the utility $u_{bj}$ is a concave function with respect to $x_{bj}$. As the budget allocation to battlefield $j$ increases, its marginal utility decreases until it matches the marginal utility of other battlefields with initially lower marginal utility. From this point, the follower distributes their budget across these battlefields, maintaining equal marginal utility, until it eventually decreases to an even lower level. Ultimately, the follower exhausts their entire budget. The lowest marginal utility is given by $\frac{\left(\sum_{l=1}^{k^{*}}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum_{l=1}^{k^{*}}x_{al}\right)^{2}}$, as defined in $k^{*}$ in (4).
The proofs of Lemmas 7 through 10 involve applying the follower’s best response strategy. It is essential to consider whether these game-splitting and merging operations, as well as the leader’s strategies $\bm{x}_{a}$, alter the structure of the set $K(\bm{x}_{a})$. This set, $K(\bm{x}_{a})$, represents the battlefields on which the follower allocates positive budgets. Essentially, there is a mapping between the leader’s commitment set and the follower’s strategy set $\bm{x}_{b}$ across the original game $\mathcal{G}$ and the subsequent games $\mathcal{G}^{(k)}$ for $k=1,2,3$.

### v3

@@ Abstract
Resource competition problems are often modeled using Colonel Blotto games, where players take simultaneous actions. However, many real-world scenarios involve sequential decision-making rather than simultaneous moves.
To model these dynamics, we represent the Lottery Colonel Blotto game as a Stackelberg game, in which one player, the leader, commits to a strategy first, and the other player, the follower, responds. We derive the Stackelberg equilibrium for this game, formulating the leader’s strategy as a bi-level optimization problem.
To solve this, we develop a constructive method based on iterative game reductions, which allows us to efficiently compute the leader’s optimal commitment strategy in polynomial time. Additionally, we identify the conditions under which the Stackelberg equilibrium coincides with the Nash equilibrium. Specifically, this occurs when the budget ratio between the leader and the follower equals a certain threshold, which we can calculate in closed form. In some instances, we observe that when the leader’s budget exceeds this threshold, both players achieve higher utilities in the Stackelberg equilibrium compared to the Nash equilibrium. Lastly, we show that, in the best case, the leader can achieve an infinite utility improvement by making an optimal first move compared to the Nash equilibrium.
@@ 1 Introduction
Competing for resources with limited budgets in strategic settings has diverse and impactful applications. These scenarios arise in areas such as electoral competition, security, crowdsourcing and recommendation systems (Behnezhad et al., 2018; Pagey et al., 2023; Haggiag et al., 2022; Anwar et al., 2024). Many of these applications are modeled by Colonel Blotto games (Borel, 1921), which have garnered significant attention from researchers across disciplines, including computer science, economics, and sociology (Kohli et al., 2012; Jackson and Nei, 2015; Fu and Iyer, 2019).
@@
Colonel Blotto games have numerous variants, where players allocate limited resources across multiple battlefields. Resources can be discrete, like troops, or continuous, like bid prices in auctions. Player utility on each battlefield depends on the invested resources and the battlefield’s outcome function, which can follow either a winner-takes-all rule or a proportional rule. The latter, known as the Lottery Colonel Blotto game, is the focus of this paper.
@@
Colonel Blotto games are highly versatile, and most existing research treats them as normal-form games, emphasizing equilibrium existence and computation in restricted settings (Roberson, 2006; Hortala-Vallve and Llorente-Saguer, 2012; Macdonell and Mastronardi, 2015; Perchet et al., 2022). However, many real-world scenarios involve sequential decision-making rather than simultaneous play. For example, in marketplaces such as advertising auctions, e-commerce, and cloud services, large enterprises often act as leaders, committing to budget distribution strategies that smaller retailers observe and respond to. These dynamics highlight the importance of studying Colonel Blotto games in sequential settings, like Stackelberg games, where one player commits to a strategy before the other makes its decision (Hicks, 1935; Schelling, 1960). In this paper, we model the Lottery Colonel Blotto game as a Stackelberg game.
@@
The Stackelberg game can be viewed as a sequential game. Many previous studies have focused on scenarios involving an attacker as the follower and a defender as the leader, both operating under cost constraints rather than budget limitations. In these models, each player determines their effort to maximize their individual payoffs (Zhuang and Bier, 2007; Cavusoglu et al., 2008; Hausken et al., 2008; Hausken, 2012). However, Iliaev et al. (2023) point out that models without budget constraints are easier to solve for Nash equilibria, as they reduce to a single-battlefield model with multiple values. Iliaev et al. (2023) also consider budget constraints in their comparison of sequential and simultaneous games. However, they only impose budget constraints on the follower, while the first mover incurs costs without any budget limitations, and both players have symmetric valuations for the battlefield. In contrast, we propose a more general model for broader applicability, in which both the leader and the follower have budgets, and the battlefield values are asymmetric. We also compare the sequential game (the Stackelberg equilibrium) with the simultaneous game (Nash equilibria) to explore the differences in outcomes under these conditions.
@@
To compare sequential and simultaneous games, we analyze the Stackelberg equilibrium and Nash equilibria. However, obtaining a closed-form representation of Nash equilibria in Colonel Blotto games is a known challenge (Perchet et al., 2022; Li and Zheng, 2022). Such representations have only been derived in specific cases, such as when battlefield values are identical (Roberson, 2006) or when players have symmetrical budgets (Hortala-Vallve and Llorente-Saguer, 2012; Boix-Adserà et al., 2020). We focus on comparing the Stackelberg equilibrium with Nash equilibria, particularly in cases where Nash equilibria can be explicitly solved. In these solvable instances, we have made some interesting observations. For the Stackelberg equilibrium, the key challenge is computing the leader’s optimal commitment. Previous work has suggested negative conclusions about this problem. Specifically, finding an optimal pure strategy for the leader is NP-hard in normal-form games (Conitzer and Sandholm, 2006; Korzhyk et al., 2010; Letchford and Conitzer, 2010), as it involves solving a bi-level optimization problem (Renou, 2009). To address this, we explore new insights into optimal commitment and best-response dynamics, which allow us to compare the Stackelberg equilibrium with Nash equilibria in these games.
@@ 1.1 Our Contribution
  • By reformulating the leader’s objective and reducing the search space for the follower’s best response strategy, we can compute the leader’s optimal commitment strategy by a polynomial number of iterations.
  • We provide the sufficient and necessary conditions, under which the Stackelberg equilibrium coincides with the Nash equilibria of the simultaneous-move game.
  • Finally, we address our motivating questions by constructing extreme cases that compare the leader’s utility when using the optimal commitment strategy to its utility in Nash equilibria. We observe that this ratio is also dependent on the leader’s budget relative to the follower’s budget. Furthermore, we provide an example illustrating how the follower’s utility may still increase when faced with the leader’s optimal commitment.
@@ 1.2 Related Work
When the outcome function is proportional, the probability of a player winning a battlefield depends on the proportion of resources they allocate compared to the total resources allocated by all players. This type of Colonel Blotto games is also called the Lottery Colonel Blotto game. In this line of work, it is called symmetric if all players have equal budgets; otherwise, it is asymmetric. When all battlefields have the same value, and this value is consistent for all players, the game is homogeneous. If the battlefields possess different values, but these values remain the same for every player, the game is heterogeneous. Finally, when the battlefields’ values differ, and these values vary for each player, the game is termed a generalized game. The Lottery Colonel Blotto game has been proven to possess pure Nash equilibria (Kim et al., 2018). Friedman (1958) shows that the pure Nash equilibrium is unique and exhibits proportionality features in a two-player setup with heterogeneous battlefield values and asymmetric player budgets. Duffy and Matros (2015) extend Friedman’s analysis from two players to multiple players. Kim et al. (2018) provide a method to identify all pure Nash equilibria in the case of two players with asymmetric budgets and multiple battlefields with generalized values, and present an example where the pure Nash equilibrium is not unique. Kovenock and Arjona (2019) provide the best response given the strategy of the other player. However, none of the existing work has provided a closed-form representation of the pure Nash equilibria (Kim et al., 2018; Kovenock and Arjona, 2019). In a variant where the outcome function of a battlefield is determined by the Tullock rent-seeking contest success function, the proportional rule is parameterized by a battle-specific discriminatory power and a battle-and-contestant-specific lobbying effectiveness (Xu and Zhou, 2018; Osorio, 2013; Li and Zheng, 2022). In this variant, the existence of a pure Nash equilibrium is not always guaranteed. Xu et al. (2022) provide sufficient conditions for the existence of the pure Nash equilibria, while Li and Zheng (2022) analyze the properties of the pure Nash equilibria and provide sufficient conditions for its uniqueness.
@@ 2 Preliminaries
$u_{i}(\bm{x}_{a},\bm{x}_{b})=\sum_{j=1}^{n}u_{ij}=\sum_{j=1}^{n}\frac{x_{ij}\cdot v_{ij}}{x_{ij}+x_{-ij}},\;\forall i\in\{a,b\}.$ (2)
When $x_{aj}>0$, for all $j\in[n]$, Kovenock and Arjona (2019) characterize the other player’s best response function as follows.
@@
Kovenock and Arjona (2019) Given the leader’s strategy $\bm{x}_{a}=(x_{aj})_{j=1}^{n}$, where $x_{aj}>0$, $\forall j\in[n]$, assume without loss of generality that all battlefields are ordered such that $\frac{v_{b1}}{x_{a1}}\geq\frac{v_{b2}}{x_{a2}}\geq\cdots\geq\frac{v_{bn}}{x_{an}}$. The unique optimal budget allocation of the follower, $\bm{x}_{b}=(x_{b1},x_{b2},\cdots,x_{bn})$, to battlefield $j\in[n]$ is characterized as follows:
$x_{bj}=\begin{cases}\frac{\left(x_{aj}v_{bj}\right)^{\frac{1}{2}}\left(x_{b}+\sum\limits_{j^{\prime}\in K(\bm{x}_{a})}x_{aj^{\prime}}\right)}{\sum\limits_{j^{\prime}\in K(\bm{x}_{a})}\left(x_{aj^{\prime}}v_{bj^{\prime}}\right)^{\frac{1}{2}}}-x_{aj},&\text{if }j\in K(\bm{x}_{a});\\ 0,&\text{if }j\in[n]\backslash K(\bm{x}_{a}),\end{cases}$ (3)
@@
$k^{*}=\max\left\{k\in[n]:\frac{v_{bj}}{x_{aj}}>\frac{\left(\sum\limits_{l=1}^{j}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum\limits_{l=1}^{j}x_{al}\right)^{2}},\forall j\leq k\right\}.$ (4)
This characterization is a useful tool for our subsequent analysis. To use this result to characterize the leader’s optimal commitment, we first assume that the leader will allocate a positive budget to every battlefield. Following a series of constructions in Section 3, we identify the support of the follower’s best response strategies and provide a closed-form expression of the leader’s optimal commitment strategy in Section 4. To complete this analysis, we will verify that the leader indeed allocates a positive budget to every battlefield in the optimal commitment strategy. We focus on the pure strategy of the leader.
@@ 3 The Support of the Follower’s Best Response Strategy
    Consider a game $\mathcal{G}$, with the leader’s commitment strategy denoted by $\bm{x}_{a}$ (which may not be optimal), and the follower’s best response denoted by $\bm{x}_{b}$. If we divide a single battlefield into multiple sub-battlefields in such a way that the leader and follower’s valuations of the original battlefield are evenly distributed among these sub-battlefields, the resulting game is denoted by $\mathcal{G}^{(1)}$. By evenly distributing their budgets across these sub-battlefields, the utilities for both the leader and the follower remain unchanged in $\mathcal{G}^{(1)}$. The formal statement and its proof are deferred to Lemma 6.
@@
    By repeatedly performing this battlefield-splitting process, we create a game in which the follower values all battlefields equally. Without loss of generality, we can rename the battlefields so that the leader values them in increasing order. Denote this game by $\mathcal{G}^{(2)}$. According to Lemma 7, in the optimal commitment strategy, the leader should allocate more or at least equal budget to battlefields that they value strictly higher in order to achieve a strictly greater payoff.
@@
    Furthermore, we observe that, in the optimal commitment strategy, when the leader values two battlefields equally, they should allocate an equal budget to both battlefields. For a formal statement and proof, see Lemma 8.
@@
    Following the above procedure, we obtain a game where $v_{b1}=v_{b2}=\dots$, meaning the follower values all battlefields equally. Meanwhile, the leader values the battlefields in increasing order, with some values strictly increasing and others remaining equal. Our final operation is to merge these battlefields that split from the same battlefield. In this new game $\mathcal{G}^{(3)}$, the leader and follower’s values and budgets across these sub-battlefields are merged into the original battlefield. Their utilities remain unchanged from $\mathcal{G}^{(2)}$. The formal statement and proof are deferred to Lemma 9.
With these constructions and observations, we now provide the main result of this section.
@@
$\frac{\partial u_{bj}}{\partial x_{bj}}=\frac{\partial(\frac{x_{bj}v_{bj}}{x_{aj}+x_{bj}})}{\partial x_{bj}}=\frac{x_{aj}v_{bj}}{(x_{aj}+x_{bj})^{2}},$ (5)
$\frac{\partial^{2}u_{bj}}{\partial x_{bj}^{2}}=\frac{-2x_{aj}v_{bj}}{(x_{aj}+x_{bj})^{3}}<0.$
Hence, the utility $u_{bj}$ is a concave function with respect to $x_{bj}$. As the budget allocation to battlefield $j$ increases, its marginal utility decreases until it matches the marginal utility of other battlefields with initially lower marginal utility. From this point, the follower distributes its budget across these battlefields, maintaining equal marginal utility, until it eventually decreases to an even lower level. Ultimately, the follower exhausts its entire budget. The lowest marginal utility is given by $\frac{\left(\sum_{l=1}^{k^{*}}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum_{l=1}^{k^{*}}x_{al}\right)^{2}}$, as defined in $k^{*}$ in (4).
It is essential to consider whether these game-splitting and merging operations, as well as the leader’s strategies $\bm{x}_{a}$, alter the structure of the set $K(\bm{x}_{a})$. This set, $K(\bm{x}_{a})$, represents the battlefields on which the follower allocates positive budgets. Essentially, there is a mapping between the leader’s commitment set and the follower’s strategy set $\bm{x}_{b}$ across the original game $\mathcal{G}$ and the subsequent games $\mathcal{G}^{(k)}$ for $k=1,2,3$.